OpenStudy (anonymous):

From a set of first 10 natural numbers,three distinct prime numbers a,b,c are selected to form a quadratic equation ax^2+bx+c=0 having real roots.find the sum of the roots of all such possible quadratic equations that can be formed.

OpenStudy (anonymous):

too tired

OpenStudy (anonymous):

please try it..!if u can

OpenStudy (anonymous):

is 0 included in "natural numbers"

OpenStudy (anonymous):

possible prime numbers are 2, 3, 5, 7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to find the roots? If so, use all possible combinations of 2,3,5,7 for a,b, and c , and find roots of all equations. add the roots of the equations with real roots together.

OpenStudy (anonymous):

You can exclude some values of b 2 and 3 will never work out to real roots

OpenStudy (anonymous):

Also for b = 5 and b = 7 only values or 2,3 for a and c will work. 4 quadratics

OpenStudy (anonymous):

for roots to be real D >0.

OpenStudy (anonymous):

that is b^2 - 4ac > 0

OpenStudy (anonymous):

total 5 quadratic equations i think can be possible where the sum of roots will be = -b/a

OpenStudy (anonymous):

am I correct?

OpenStudy (anonymous):

I only see 4 myself.

OpenStudy (anonymous):

2x^2 + 5x + 3, 3x^2 + 5x + 2, 2x^2 + 7x + 3, 3x^2 + 7x + 2

OpenStudy (anonymous):

Giving roots of -5/4+-1, -5/6+-1, -7/4+-5, -7/6+-5

OpenStudy (anonymous):

when b=5 , a can take 2 or 3 when b=7 a can take 2,5,3

OpenStudy (anonymous):

You're right, 2 and 5 work as a and c with 7. That gives two more equations.

OpenStudy (anonymous):

-7/4+-3 and -7/6+-3

OpenStudy (anonymous):

How come you arrive to a conclusion of -5/4+-1 as there roots As I know sum of the roots will be -b/a so it will be -5/2,-5/3,-7/2,-7/5,-7/3

OpenStudy (anonymous):

That is my bad notation I mean -5/4(+-)1 as in -5/4 + 1 and -5/4 - 1

OpenStudy (anonymous):

-5/4 + 1 and -5/4 - 1 ???? can you please explain?

OpenStudy (anonymous):

That would come out -b/a for the sum of both.

OpenStudy (anonymous):

ok. Consider one of the quadratic equation being 2x^2+5x+3 The sum of roots in this case will be -5/2

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

similarly you have 2x^2+5x+3, 3x^2+5x+2, 2x^2+7x+3, 2x^2+7x+5, 3x^2+7x+2, 3x^2+7x+5, 5x^2+7x+2, 5x^2+7x+3,

OpenStudy (anonymous):

So you'll have what you wrote above and -7/2 again since 2x^2 + 7x + 3 and 2x^2 + 7x + 5 both give -7/2

OpenStudy (anonymous):

49 < 4*5*3

OpenStudy (anonymous):

So not 5x^2 + 7x + 3 or 3x^2 + 7x + 5

OpenStudy (anonymous):

Otherwise you have them all.

OpenStudy (anonymous):

yes in that case D< 0 and roots will be imaginary

OpenStudy (anonymous):

So finally we have 2x^2+5x+3, 3x^2+5x+2, 2x^2+7x+3, 2x^2+7x+5, 3x^2+7x+2,5x^2+7x+2

OpenStudy (anonymous):

That is it.

OpenStudy (anonymous):

I get 149/10 out of mathematica as the final sum

OpenStudy (anonymous):

so it will be -5/2-5/3-7/2-7/2-7/3-7/5 = -149/10

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Correct...Thanks!!!!I appreciate it.

OpenStudy (anonymous):

No problem